Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2017-03-05

Observers: Anders Nyholm, Cristina Barbarino
TAT on duty: Joe Anderson, Lluís Galbany
Support team: Francesco Taddia, Christoffer Fremling, Jesper Sollerman

Daytime calibrations: PESSTO_calib, PESSTO_bias
Evening skyflats: UVi, BR
Morning calibrations: PESSTO_calib-onlyarcs

Weather:  Clear skies. Humidity around 40 % most of the night, around 30 % in the morning.

Seeing: 0.8 arcsec (23:35 UT), 0.5 arcsec (02:45 UT), 0.8 arcsec (05:10 UT), 0.5 arcsec (09:10 UT)

UT/Target/RA/DEC/time spent(min)/slit/grism or filter /description
00:12 SN2016iae 04:12:05.53 -32:51:44.7 60  Follow-up, gr11, strong signal from SN2016iae (SN2016ija in same slit, weak signal)
01:02 Image analysis
01:17 ATLAS17bds 08:56:46.74 +29:55:15.5 37  Follow-up, gr11, at high airmass, strong trace. Blue continuum with some features. SNID match (at z = 0.033) to a SN IIL?
01:54 AT2017aww 08:37:38.21 +19:24:13.9 38  Follow-up, gr13, strong trace. Spectrum with emission features. Double-peaked H alpha?
02:33 AT2016jbu 07:36:25.96 -69:32:55.2 60  Follow-up, gr11 and gr16 (strong trace)
03:39 L745a 07:40:20.79 -17:24:49.1 16  Standard, gr11+gr13+gr16
03:55        Problem with the EFOSC ICS (Instrument Control Software). TO rebooting.
04:24        TO goes to NTT to manually inspect the problem. Mechanical problem, eventually fixed.
04:47 PS17bek 10:47:41.90 +26:50:06.0 109  Follow-up, gr11 (clear trace), gr16 (weaker trace), at high airmass
06:39 ATLAS17bas 11:07:35.49 +22:51:04.4 60  Follow-up, gr11+gr16, at high airmass
07:45 SN2017yv 10:23:40.39 -35:49:30.7 60  Follow-up, gr11+gr16 (clear traces)
08:53 AT2017bkc 17:50:29.96 -01:48:09.0 60  Follow-up, gr11+gr16 (clear traces)
09:59  EG274 16:23:33.75 -39:13:47.5 13  Standard, gr11+gr13+gr16
Comments